OpenStudy (anonymous):

In a triangle, the measure of the first angle is 24 degrees more than the measure of the second angle. The measure of the third angle is four times the measure of the second angle. If the total number of degrees in a triangle is 180, what is the measure of the largest angle?

Parth (parthkohli):

Hi, we can use variables here!

Parth (parthkohli):

OK. Let's say that the measure of the first angle is \(x\). What is the measure of the second angle in terms of \(x\)?

OpenStudy (anonymous):

ok in the problem it says ur first angle is 24+x

OpenStudy (anonymous):

ur second angle is 4x

OpenStudy (anonymous):

ur third angle is x

Parth (parthkohli):

if the first angle is 24 + x, then the second angle should be considered x.

OpenStudy (anonymous):

the total of all the angles is 180

Parth (parthkohli):

x, x - 24 and 4(x-24) are the angles where x is the first angle.

OpenStudy (anonymous):

therefor (24+x)+(4x)+(x)=180

OpenStudy (anonymous):

subtract the 24 so x+4x+x=156

OpenStudy (anonymous):

factor out the x

OpenStudy (anonymous):

x(1+4+1)=156

OpenStudy (anonymous):

x(6)=156 then divide the 6

OpenStudy (anonymous):

x=26

OpenStudy (anonymous):

plug in x into the three orginal formuals and u shoud get 26, 50, 104 which will give u 180

OpenStudy (anonymous):

sorry for all the steps lol gl